# Scene123: One Prompt to 3D Scene Generation via Video-Assisted and Consistency-Enhanced MAE

### v1

@@ Abstract
As Artificial Intelligence Generated Content (AIGC) advances, a variety of methods have been developed to generate text, images, videos, and 3D objects from single or multimodal inputs, contributing efforts to emulate human-like cognitive content creation. However, generating realistic large-scale scenes from a single input presents a challenge due to the complexities involved in ensuring consistency across extrapolated views generated by models. Benefiting from recent video generation models and implicit neural representations, we propose Scene123, a 3D scene generation model, that not only ensures realism and diversity through the video generation framework but also uses implicit neural fields combined with Masked Autoencoders (MAE) to effectively ensures the consistency of unseen areas across views. Specifically, we initially warp the input image (or an image generated from text) to simulate adjacent views, filling the invisible areas with the MAE model. However, these filled images usually fail to maintain view consistency, thus we utilize the produced views to optimize a neural radiance field, enhancing geometric consistency. Moreover, to further enhance the details and texture fidelity of generated views, we employ a GAN-based Loss against images derived from the input image through the video generation model. Extensive experiments demonstrate that our method can generate realistic and consistent scenes from a single prompt. Both qualitative and quantitative results indicate that our approach surpasses existing state-of-the-art methods. We show encourage video examples at https://yiyingyang12.github.io/Scene123.github.io/.
Figure 1: Some examples generated by Scene123. For a single input image or text, our method can generate 3D scenes with consistent perspectives, fine geometry, and realistic textures, applicable to real, virtual, or object-centered scenes.
@@ 1 Introduction
3D scene generation aims to create realistic or stylistically specific scenes from limited prompts, such as a few images or a text description. This is a fundamental issue in computer vision and graphics and a critical challenge in generative artificial intelligence. Recent advancements have demonstrated substantial progress through the use of vision-language models [58, 37], generative models such as Generative Adversarial Networks (GANs) [96], Variational autoencoders (VAEs) [67], or diffusion models [88], and scene representations like Neural Radiance Fields (NeRF) [49] or 3D Gaussian Splatting (3DGS) [36]. Typically, these methods begin by generating images with a pretrained generative model [63, 90], or by directly using images for image-to-scene generation, then estimate additional 3D surface geometric details such as depth and normal [54, 38], and subsequently render the surface textures of the scenes using generative strategies like inpainting [85, 1]. Moreover, some approaches update the geometric surfaces as new views are generated to maintain the coherence of the scene [95]. However, these methods often rely on pre-trained models [58, 37], resulting in inconsistencies and artifacts in the generated scenes. Additionally, these methods face challenges in producing high-quality, coherent 3D representations across diverse and complex environments.
@@
It is noteworthy that scene synthesis based on multi-view images is a long-studied topic from the early Multi-View Stereo (MVS)  [70, 69] to recent implicit neural representations [92, 2], ensuring view consistency in generated scenes through multi-view matching mechanisms. However, when the input is reduced to a single image or a textual description, lacking reference multi-view images, the performance of these methods is greatly compromised. Fortunately, methodologies such as Masked Autoencoders (MAE)  [23] provide avenues for extrapolating to areas unseen in new views, while the incorporation of additional semantic layers augments the coherence of the synthesized scenes. Additionally, video generation models  [26, 44] facilitate the enhancement of scenes with richer priors and more detailed texture information. Thus, how to utilize multi-view reconstruction with robust physical constraints alongside effective expansions in new viewpoints and video generation models is still an unreached area.
@@
Scene123 investigates a methodology for 3D scene reconstruction employing stringent physical constraints alongside a robust multi-view MAE and video generation models to ensure view consistency. To achieve this, for each input or generated single image, we first create images of nearby perspectives through warping. Subsequently, a consistency-enhanced MAE is designed to inpaint the unseen areas, with a shared codebook maintained to distribute global information to every invisible area. We implement a progressive strategy derived from Text2NeRF  [95] to incrementally update perspectives throughout this process. Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images. The synergistic operation of these two modules enables Scene123 to generate consistent, finely detailed three-dimensional scenes with photo-realistic textures from a single prompt.
@@
We conduct extensive experiments on text-to-scene and image-to-scene generation, encompassing both real and virtual scenes. The results offer robust empirical evidence that strongly supports the effectiveness of our framework. Our contributions can be summarized as: 1) We propose a novel scene generation framework based on one prompt, which establishes the connection between MAE and video generation models for the first time to ensure view consistency and realism of the generated scenes. 2) The Consistency-Enhanced MAE is designed to fill unseen areas in new views by injecting global semantic information and combining it with neural implicit fields, ensuring consistent surface representation across various views. 3) We introduce the video-assisted 3D-aware generative refinement module, which enhances scene reconstruction by integrating the diversity and realism of video generation models through a GAN-based function to significantly improve detail and texture fidelity. 4) Extensive experiments validate the efficacy of Scene123, demonstrating greater accuracy in surface reconstruction, higher realism in reconstructed views, and better texture fidelity compared to the state-of-the-art methods. The code and weights will be available.
@@ 2 Related Work
3D Scene Generation. Recent advancements in text-to-3D and image-to-3D scene generation have been remarkable. Most research in this domain can be classified into two main categories: object-centric [14, 34, 40, 41, 47, 56, 57, 60, 72, 74, 78, 86, 92], which focuses on objects without backgrounds, and holistic, which involves generating entire 3D scenes or 3D trajectories including backgrounds. Expanding these techniques to create fully unconstrained 3D scenes remains a significant challenge. Many recent advancements in text-driven 3D scene generation have focused on modeling 3D scenes using text inputs [32, 13, 95, 19, 28, 94, 73, 97]. Due to the scarcity of paired text-3D scene data, most studies utilize Contrastive Language-Image Pre-training (CLIP) [58] or pre-trained text-to-image models to interpret the text input. Text2Scene [32] employs CLIP to model and stylize 3D scenes from text (or image) inputs by decomposing the scene into manageable sub-parts. Set-the-Scene [13] and Text2NeRF [95] generate NeRFs from text using text-to-image diffusion models to represent 3D scenes. SceneScape [19] and Text2Room [28] leverage a pre-trained monocular depth prediction model for enhanced geometric consistency and directly generate the 3D textured mesh representation of the scene. However, these methods depend on inpainted images for scene completion, which, while producing realistic visuals, suffer from limited 3D consistency. More recent studies [20, 2, 80] have successfully generated multi-object compositional 3D scenes. Another class of methods utilizes auxiliary inputs, such as layouts [55, 17], to enhance scene generation. Unlike text captions, which can be rather vague, some approaches generate 3D scenes from image inputs, where the 3D scene corresponds closely to the depicted image. Early scene generation methods often require specific scene data for training to obtain category-specific scene generators, such as GAUDI [3], or focus on single scene reconstruction based on the input image, such as PixelSynth [61] and Worldsheet [31]. However, these methods are often limited by the quality of the generation or the extensibility of the scene.
@@
Recently, numerous studies have concentrated on generating 3D scenes from image inputs. PERF [83] generates 3D scenes from a single panoramic image, using diffusion models to supplement shadow areas. ZeroNVS [68] extends this capability by reconstructing both objects and environments in 3D from a single image. Despite its environmental reconstruction lacking some detail, the algorithm demonstrates an understanding of environmental contexts. LucidDreamer [12] and WonderJourney [93] utilize general-purpose depth estimation models to project hallucinated 2D scene extensions into 3D representations. However, these methods still face challenges in achieving realism, often producing artifacts due to the reliance on pre-trained models.
@@
Video Diffusion Models and 3D-aware GANs. Diffusion models [76] have recently emerged as powerful generative models capable of producing a wide array of images [5, 63, 65] and videos [4, 22, 81] by iteratively denoising a noise sample. Among these models, the publicly available Stable Diffusion (SD) [63] and Stable Video Diffusion (SVD) [4] exhibit strong generalization capabilities due to training on extremely large datasets such as LAION [71] and LVD [4]. Consequently, they are frequently used as foundational models for various generation tasks, including novel view synthesis. To enhance generalization and multi-view consistency, some contemporary works leverage the temporal priors in video diffusion models for object-centric 3D generation. For instance, IM-3D [46] and SV3D [82] explore the capabilities of video diffusion models in object-centric multi-view generation. V3D [11] extends this approach to scene-level novel view synthesis. However, these methods often produce unsatisfactory results for complex objects or scenes, leading to inconsistencies among multiple views or unrealistic geometries.
@@
Early works, like 3D-GAN [87], Pointflow [89], and ShapeRF [6] focus more on the category-specific texture-less geometric shape generation based on the representations of voxels or point clouds. However, limited by the generation capabilities of GANs, these methods can only generate rough 3D assets of specific categories. Subsequently, HoloGAN [51], GET3D [21], and EG3D [7] employ GAN-based 3D generators conditioned on latent vectors to produce category-specific textured 3D assets. Recently, as seen in GigaGAN [35], the Generative Adversarial Networks (GANs) methods are better suited for high-frequency details than diffusion models. Furthermore, IT3D [10] proposes a novel Diffusion-GAN dual training strategy to overcome the view inconsistency challenges. However, the training process of GAN is prone to the issue of mode collapse, which limits the diversity of generation results.
Vision Transformers. Computer vision has undergone a transformative shift with the emergence of vision transformers [16, 59]. Their ability to model global and local contexts has sparked interest in exploring methods to pre-train these backbones. MoCoV3 [9] and MAE [23] delved into different aspects of self-supervised Vision Transformers. In particular, MAE achieved state-of-the-art results by drawing inspiration from BERT [15], which involves randomly masking words in sentences, and using masked image reconstruction. Subsequent research extended this concept to the 3D domain [24, 39, 52, 77], leading to impressive results in downstream 3D vision tasks.
@@
Figure 2: Scene123’s pipeline includes two key modules: the consistency-enhanced MAE and the 3D-aware generative refinement module. The former generates adjacent views from an input image via warping, using the MAE model to inpaint unseen areas with global semantics and optimizing an implicit neural field for viewpoint consistency. The latter generates realistic scene videos from the input image with a pre-trained video generation model, enhancing realism through adversarial loss with rendered images.
@@ 3.1 Overview
In this paper, we propose a novel 3D scene generation framework based on one prompt, a single image or textual description, ensuring viewpoint consistency and realistic surface textures for both real and synthetically styled scenes, as shown in Fig. 2. We first design the Consistency-Enchanced MAE module to fill unseen areas in novel views by injecting global semantic information in Sec. 3.2. With support views in the initialized database $\mathbf{S}$ generated by the Consistency-Enhanced MAE module, we employ a NeRF network to represent the 3D scene as the physical constraints of view consistency in Sec. 3.3. Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images in  3.4. Finally, the optimization and implementation details of the model are detailed in Sec. 3.5.
@@ 3.2 Consistency-Enhanced MAE Scene Completion
Scene Initialization. Given the reference image $\mathbf{I_{0}}$, notably, for only text prompt input, we utilize the stable diffusion model [63] to generate initial image $\mathbf{I_{0}}$, we then feed this image into the off-the-shelf depth estimation model [48], and take the output as a geometric prior for the target scene, denoted as $\mathbf{D_{0}}$. Inspired by [95], we construct an original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$ via the depth image-based rendering (DIBR) method [18], where $N$ denotes the number of initial viewpoints. Specifically, for each pixel $x$ in $\mathbf{I_{0}}$ and its depth value $y$ in $\mathbf{D_{0}}$, we compute its corresponding pixel $x_{0\rightarrow m}$ and depth $y_{0\rightarrow m}$ on a surrounding view $m$, $\left[x_{0\rightarrow m},y_{0\rightarrow m}\right]^{T}=\mathbf{K}\mathbf{P}_{m}\mathbf{P}_{0}^{-1}\mathbf{K}^{-1}[x,y]^{T}$, where $\mathbf{K}$ and $\mathbf{P}_{m}$ indicate the intrinsic matrix and the camera pose in view $m$. This database provides additional views and depth information, which could prevent the model from overfitting to the initial view.
@@
MAE Scene Completion. However, the original database $\mathcal{S}_{0}$ will inevitably have missing content since the information in the initial scene is derived from the single image $\mathbf{I_{0}}$ to construct the initialized database $\mathcal{S}$. Directly applying the original database to the 3D scene representation would inevitably suffer from limited 3D consistency of the generated scenes [95, 28]. Inspired by  [23, 30, 94], we design a Consistency-Enhanced MAE module to effectively ensure the consistency of unseen areas across views. Specifically, we first design a discrete codebook to distribute global information to every invisible area in the original database. Specifically, we use a powerful pre-trained model on the ImageNet dataset [66]. We represent the codebook as =$\mathcal{E}=\{e_{1},e_{2},...,e_{N}\}\in\mathbb{R}^{N\times n_{q}}$ , where $N$ stands for the total count of prototype vectors, $n_{q}$ denotes the dimensionality of individual vectors, and $e_{i}$ symbolizes each specific embedding vector. Specifically, given the input image $x\in\mathbb{R}^{H\times W\times 3}$, the VQ-VAE [79] employs an encoder $E$ to extract a continuous feature representation: $\hat{z}=E(x)\in\mathbb{R}^{h\times w\times c}$, where $h$ and $w$ are the height and width of the feature map. This continuous feature map $\hat{z}$ is then subjected to a quantization process $Q$, aligning it with its nearest codebook entry $e_{k}$ to obtain its discrete representation $z_{q}$, $z_{q}=Q(\hat{z}):=\underset{e_{k}\in\mathcal{E}}{\operatorname{argmin}}\left\|\hat{z}_{ij}-e_{k}\right\|_{2}$, where $\hat{z}_{ij}\in\mathbb{R}^{c}$.
@@
With the pre-trained codebook, we then utilize the MAE encoder to encode the input images from the original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$, forming the image-conditional $\mathbf{s_{c}}=\{s_{1},s_{2},...,s_{M}\}$, each embedding vector $s_{i}$ queries the valuable prior information from the given codebook via the cross-attention mechanism, $\mathrm{Q}\leftarrow f_{\mathrm{q}}(\mathbf{s_{c}}),\mathrm{K}\leftarrow f_{\mathrm{k}}(\mathcal{E}),\mathrm{V}\leftarrow f_{\mathrm{v}}(\mathcal{E}),\mathbf{s_{u}}\leftarrow\text{Cross-Attention}(\mathrm{Q},\mathrm{K},\mathrm{V})=\text{Softmax}(\frac{\mathrm{Q}\mathrm{K}^{\mathrm{T}}}{\sqrt{d_{k}}})$, where $f_{q}$, $f_{k}$, and $f_{v}$ are the query, key, and value linear projections, respectively. Consequently, the global semantic information contained in the codebook is maintained to distribute global information to every invisible area. Then, we utilize the MAE decoder to decode the feature $\mathbf{s_{u}}$, deriving the initialized database $\mathcal{S}$, as shown in Fig. 2.

### v2

@@ Abstract
As Artificial Intelligence Generated Content (AIGC) advances, a variety of methods have been developed to generate text, images, videos, and 3D shapes from single or multimodal inputs, contributing efforts to emulate human-like cognitive content creation. However, generating realistic large-scale scenes from a single input presents a challenge due to the complexities involved in ensuring consistency across extrapolated views generated by models. Benefiting from recent video generation models and implicit neural representations, we propose Scene123, a 3D scene generation model, which combines a video generation framework to ensure realism and diversity with implicit neural fields integrated with Masked Autoencoders (MAE) to effectively ensure the consistency of unseen areas across views. Specifically, the input image (or a text-generated image) is first warped to simulate adjacent views, with the invisible regions filled using the consistency-enhanced MAE model. Nonetheless, the synthesized images often exhibit inconsistencies in viewpoint alignment, thus we utilize the produced views to optimize a neural radiance field, enhancing geometric consistency. Moreover, to further enhance the details and texture fidelity of generated views, we employ a GAN-based Loss against images derived from the input image through the video generation model. Extensive experiments demonstrate that our method can generate realistic and consistent scenes from a single prompt. Both qualitative and quantitative results indicate that our approach surpasses existing state-of-the-art methods. We show encourage video examples at https://yiyingyang12.github.io/Scene123.github.io/.
Figure 1: Some examples generated by our Scene123. For a single input image or text, our method can generate 3D scenes with consistent views, fine geometry, and realistic textures, applicable to real, virtual, or object-centered scenes.
@@ 1 Introduction
3D scene generation aims to create realistic or stylistically specific scenes from limited prompts, such as a few images or a text description. This is a fundamental issue in computer vision and graphics and a critical challenge in generative artificial intelligence. Recent advancements have demonstrated substantial progress through the use of vision-language models (Radford et al. 2021), generative models such as Generative Adversarial Networks (GANs) (Zhang et al. 2019), Variational autoencoders (VAEs) (Sargent et al. 2023a; Yang et al. 2023b), or diffusion models (Yang et al. 2023a), and scene representations like Neural Radiance Fields (NeRF) (Mildenhall et al. 2021; Yin et al. 2022; Ding et al. 2023; Lu et al. 2023; Yang et al. 2024) or 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023). Typically, these methods begin by generating images with a pretrained generative model (Rombach et al. 2022), or by directly using images for image-to-scene generation, then estimate additional 3D surface geometric details such as depth and normal (Piccinelli, Sakaridis, and Yu 2023; Li et al. 2023; Yin et al. 2023; Yin and Zhou 2020), and subsequently render the surface textures of the scenes using generative strategies like inpainting (Wang et al. 2023b). Moreover, some approaches update the geometric surfaces as new views are generated to maintain the coherence of the scene (Zhang et al. 2024b). However, these methods often rely on pre-trained models (Radford et al. 2021; Li et al. 2022), resulting in inconsistencies and artifacts in the generated scenes. Additionally, these methods face challenges in producing high-quality, coherent 3D representations across diverse and complex environments.
@@
It is noteworthy that scene synthesis based on multi-view images is a long-studied topic from the early Multi-View Stereo (MVS)  (Schönberger et al. 2016; Schönberger and Frahm 2016) to recent implicit neural representations (Yu et al. 2021), ensuring view consistency in generated scenes through multi-view matching mechanisms. However, when the input is reduced to a single image or a textual description, lacking reference multi-view images, the performance of these methods is greatly compromised. Fortunately, methodologies such as Masked Autoencoders (MAE)  (He et al. 2022) provide avenues for extrapolating to areas unseen in new views, while the incorporation of additional semantic layers augments the coherence of the synthesized scenes. Additionally, video generation models  (Luo et al. 2023) facilitate the enhancement of scenes with richer priors and more detailed texture information. Thus, how to utilize multi-view reconstruction with robust physical constraints alongside effective expansions in new viewpoints and video generation models is still an unreached area.
@@
Scene123 investigates a methodology for 3D scene reconstruction employing stringent physical constraints alongside a robust multi-view MAE and video generation models to ensure view consistency. To achieve this, for each input or generated single image, we first create images of nearby perspectives through warping. Subsequently, a consistency-enhanced MAE is designed to inpaint the unseen areas, with a shared codebook maintained to distribute global information to every invisible area. We implement a progressive strategy derived from Text2NeRF  (Zhang et al. 2024b) to incrementally update perspectives throughout this process. Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images. The synergistic operation of these two modules enables Scene123 to generate consistent, finely detailed three-dimensional scenes with photo-realistic textures from a single prompt.
@@
We conduct extensive experiments on text-to-scene and image-to-scene generation, encompassing both real and virtual scenes. The results offer robust empirical evidence that strongly supports the effectiveness of our framework. Our contributions can be summarized as: 1) We propose a novel scene generation framework based on one prompt, which establishes the connection between MAE and video generation models for the first time to ensure view consistency and realism of the generated scenes. 2) The Consistency-Enhanced MAE is designed to fill unseen areas in new views by injecting global semantic information and combining it with neural implicit fields, ensuring consistent surface representation across various views. 3) We introduce the video-assisted 3D-aware generative refinement module, which enhances scene reconstruction by integrating the diversity and realism of video generation models through a GAN-based function to significantly improve detail and texture fidelity. 4) Extensive experiments validate the efficacy of Scene123, demonstrating greater accuracy in surface reconstruction, higher realism in reconstructed views, and better texture fidelity compared to the SOTA methods. The data and code will be available.
@@ 2 Related Work
Text to 3D Scene Generation. Many recent advancements in text-driven 3D scene generation have focused on modeling 3D scenes using text inputs (Hwang, Kim, and Kim 2023; Zhang et al. 2024b). Due to the scarcity of paired text-3D scene data, most studies utilize Contrastive Language-Image Pre-training (CLIP) (Radford et al. 2021) or pre-trained text-to-image models to interpret the text input. Text2Scene (Hwang, Kim, and Kim 2023) employs CLIP to model and stylize 3D scenes from text (or image) inputs by decomposing the scene into manageable sub-parts. Set-the-Scene (Cohen-Bar et al. 2023) and Text2NeRF (Zhang et al. 2024b) generate NeRFs from text using text-to-image diffusion models to represent 3D scenes. SceneScape (Fridman et al. 2024) and Text2Room (Höllein et al. 2023) leverage a pre-trained monocular depth prediction model for enhanced geometric consistency and directly generate the 3D textured mesh representation of the scene. However, these methods depend on inpainted images for scene completion, which, while producing realistic visuals, suffer from limited 3D consistency. More recent studies (Bai et al. 2023) have successfully generated multi-object compositional 3D scenes. Another class of methods utilizes auxiliary inputs, such as layouts (Po and Wetzstein 2023), to enhance scene generation. Unlike text captions, which can be rather vague, some approaches generate 3D scenes from image inputs, where the 3D scene corresponds closely to the depicted image. Early scene generation methods often require specific scene data for training to obtain category-specific scene generators, such as GAUDI (Bautista et al. 2022), or focus on single scene reconstruction based on the input image, such as PixelSynth (Rockwell, Fouhey, and Johnson 2021) and Worldsheet (Hu et al. 2021). However, these methods are often limited by the quality of the generation or the extensibility of the scene.
@@
Image to 3D Scene Generation. Recently, numerous studies have concentrated on generating 3D scenes from image inputs. PERF (Wang et al. 2023a) generates 3D scenes from a single panoramic image, using diffusion models to supplement shadow areas. ZeroNVS (Sargent et al. 2023b) extends this capability by reconstructing both objects and environments in 3D from a single image. Despite its environmental reconstruction lacking some detail, the algorithm demonstrates an understanding of environmental contexts. LucidDreamer (Chung et al. 2023) and WonderJourney (Yu et al. 2023) utilize general-purpose depth estimation models to project hallucinated 2D scene extensions into 3D representations. However, these methods still face challenges in achieving realism, often producing artifacts due to the reliance on pre-trained models.
@@
Video Diffusion Models and 3D-aware GANs. Diffusion models (Song et al. 2020) have recently emerged as powerful generative models capable of producing a wide array of images (Blattmann et al. 2023b) and videos (Blattmann et al. 2023a) by iteratively denoising a noise sample. Among these models, the publicly available Stable Diffusion (SD) (Rombach et al. 2022) and Stable Video Diffusion (SVD) (Blattmann et al. 2023a) exhibit strong generalization capabilities due to training on extremely large datasets such as LAION (Schuhmann et al. 2022) and LVD (Blattmann et al. 2023a). Consequently, they are frequently used as foundational models for various generation tasks, including novel view synthesis. To enhance generalization and multi-view consistency, some contemporary works leverage the temporal priors in video diffusion models for object-centric 3D generation. For instance, IM-3D (Melas-Kyriazi et al. 2024) and SV3D (Voleti et al. 2024) explore the capabilities of video diffusion models in object-centric multi-view generation. V3D (Chen et al. 2024b) extends this approach to scene-level novel view synthesis. However, these methods often produce unsatisfactory results for complex objects or scenes, leading to inconsistencies among multiple views or unrealistic geometries.
@@
Early works, like 3D-GAN (Wu et al. 2016), Pointflow (Yang et al. 2019), and ShapeRF (Cai et al. 2020) focus more on the category-specific texture-less geometric shape generation based on the representations of voxels or point clouds. However, limited by the generation capabilities of GANs, these methods can only generate rough 3D assets of specific categories. Subsequently, HoloGAN (Nguyen-Phuoc et al. 2019), GET3D (Gao et al. 2022), and EG3D (Chan et al. 2022) employ GAN-based 3D generators conditioned on latent vectors to produce category-specific textured 3D assets. Recently, as seen in GigaGAN (Kang et al. 2023), the Generative Adversarial Networks (GANs) methods are better suited for high-frequency details than diffusion models. Furthermore, IT3D (Chen et al. 2024a) proposes a novel Diffusion-GAN dual training strategy to overcome the view inconsistency challenges. However, the training process of GAN is prone to the issue of mode collapse, which limits the diversity of generation results.
@@
Figure 2: Scene123’s pipeline includes two key modules: the consistency-enhanced MAE and the 3D-aware generative refinement module. The former generates adjacent views from an input image via warping, using the MAE model to inpaint unseen areas with global semantics and optimizing an implicit neural field for viewpoint consistency. The latter generates realistic videos from the input image with a pre-trained video generation model, enhancing realism through adversarial loss with rendered images.
@@ 3.1 Overview
In this paper, we propose a novel 3D scene generation framework based on one prompt, a single image or textual description, ensuring viewpoint consistency and realistic surface textures for both real and synthetically styled scenes, as shown in Fig. 2. We first design the Consistency-Enchanced MAE module to fill unseen areas in novel views by injecting global semantic information. With support views in the initialized database $\mathbf{S}$ generated by the Consistency-Enhanced MAE module, we employ a NeRF network to represent the 3D scene as the physical constraints of view consistency. Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images. Finally, the optimization and implementation details of the model are detailed.
@@ 3.2 Consistency-Enhanced MAE Scene Completion
Scene Initialization. Given the reference image $\mathbf{I_{0}}$, notably, for only text prompt input, we utilize the stable diffusion model (Rombach et al. 2022) to generate initial image $\mathbf{I_{0}}$, we then feed this image into the off-the-shelf depth estimation model (Miangoleh et al. 2021), and take the output as a geometric prior for the target scene, denoted as $\mathbf{D_{0}}$. Inspired by (Zhang et al. 2024b), we construct an original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$ via the depth image-based rendering (DIBR) method (Fehn 2004), where $N$ denotes the number of initial viewpoints. Specifically, for each pixel $x$ in $\mathbf{I_{0}}$ and its depth value $y$ in $\mathbf{D_{0}}$, we compute its corresponding pixel $x_{0\rightarrow m}$ and depth $y_{0\rightarrow m}$ on a surrounding view $m$, $\left[x_{0\rightarrow m},y_{0\rightarrow m}\right]^{T}=\mathbf{K}\mathbf{P}_{m}\mathbf{P}_{0}^{-1}\mathbf{K}^{-1}[x,y]^{T}$, where $\mathbf{K}$ and $\mathbf{P}_{m}$ indicate the intrinsic matrix and the camera pose in view $m$. This database provides additional views and depth information, which could prevent the model from overfitting to the initial view.
@@
MAE Scene Completion. However, the original database $\mathcal{S}_{0}$ will inevitably have missing content since the information in the initial scene is derived from the single image $\mathbf{I_{0}}$ to construct the initialized database $\mathcal{S}$. Directly applying the original database to the 3D scene representation would inevitably suffer from limited 3D consistency of the generated scenes (Zhang et al. 2024b; Höllein et al. 2023). Inspired by  (He et al. 2022; Hu et al. 2023; Zhang et al. 2024a), we design a Consistency-Enhanced MAE module to effectively ensure the consistency of unseen areas across views. Specifically, we first design a discrete codebook to distribute global information to every invisible area in the original database. We represent the codebook as =$\mathcal{E}=\{e_{1},e_{2},...,e_{N}\}\in\mathbb{R}^{N\times n_{q}}$ , where $N$ stands for the total count of prototype vectors, $n_{q}$ denotes the dimensionality of individual vectors, and $e_{i}$ symbolizes each specific embedding vector. Specifically, given the input image $x\in\mathbb{R}^{H\times W\times 3}$, the VQ-VAE (Van Den Oord, Vinyals et al. 2017) employs an encoder $E$ to extract a continuous feature representation: $\hat{z}=E(x)\in\mathbb{R}^{h\times w\times c}$, where $h$ and $w$ are the height and width of the feature map. This continuous feature map $\hat{z}$ is then subjected to a quantization process $Q$, aligning it with its nearest codebook entry $e_{k}$ to obtain its discrete representation $z_{q}$ as follows:
$z_{q}=Q(\hat{z}):=\underset{e_{k}\in\mathcal{E}}{\operatorname{argmin}}\left\|\hat{z}_{ij}-e_{k}\right\|_{2},$ (1)
@@
where $\hat{z}_{ij}\in\mathbb{R}^{c}$. We pre-trained VQVAE on the ImageNet (Russakovsky et al. 2015) dataset to equip the codebook with a more representative and generalized feature set. Subsequently, we fine-tuned it on specific scenes to better capture and represent the unique characteristics of each scene. Moreover, we then utilize the MAE encoder to encode the input images from the original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$, forming the image-conditional $\mathbf{s_{c}}=\{s_{1},s_{2},...,s_{M}\}$, each embedding vector $s_{i}$ queries the valuable prior information from the given codebook via the cross-attention mechanism,
$\begin{split}\mathrm{Q}\leftarrow f_{\mathrm{q}}(\mathbf{s_{c}}),\qquad\enspace\mathrm{K}\leftarrow f_{\mathrm{k}}(\mathcal{E}),\qquad\enspace\mathrm{V}\leftarrow f_{\mathrm{v}}(\mathcal{E})\\ \mathbf{s_{u}}\leftarrow\text{Cross-Attention}(\mathrm{Q},\mathrm{K},\mathrm{V})=\text{Softmax}(\frac{\mathrm{Q}\mathrm{K}^{\mathrm{T}}}{\sqrt{d_{k}}}),\end{split}$ (2)
where $f_{q}$, $f_{k}$, and $f_{v}$ are the query, key, and value linear projections, respectively. Consequently, the global semantic information contained in the codebook is maintained to distribute global information to every invisible area. Then, we utilize the MAE decoder to decode the feature $\mathbf{s_{u}}$, deriving the initialized database $\mathcal{S}$, as shown in Fig. 2. Using global semantics as Key and Value in cross-attention allows the model to integrate comprehensive scene information, maintaining coherence. This approach enhances multi-view consistency by providing a unified understanding of the scene, reducing artifacts, and ensuring seamless integration of novel views.